EXERCISE 12

 The cannon with mass m = 450 kg shoots horizontally. The mass of the projectile is m = 5 kg and the speed it takes is v = 450 m / s.

 How fast will the ball bounce during the shooting?

 a) 4 m / s b) 3m / s c) 5m / s d) 2m / s


1
Expert's answer
2021-03-10T17:20:00-0500

According to the momentum conservation law:


"m_cv_c = m_pv_p"

where "m_c = 450kg, m_p = 5kg" are the masses of the cannon and the projectile respectively, and "v_c , v_p = 450m\/s" are their speeds respectively. Expressing "v_c", obtain:


"v_c = \dfrac{m_pv_p}{m_c}\\
v_c = \dfrac{5kg\cdot 450m/s}{450kg} = 5m/s"

Answer. 5 m/s.